# In-depth characterization of the Kepler-10 three-planet system with HARPS-N RVs and Kepler TTVs

### v1

@@ 2.2.2 HIRES Radial velocities
The HIRES Kepler-10 spectra were recorded using different deckers defining the length of the effective slit. The first 17 HIRES observations spanning about 38 days were made with the B5-decker (0.861 x 3.5 $\rm arcsec^{2}$) [^2] https://www2.keck.hawaii.edu/inst/hires/manual2.pdf. The longer C2-decker (0.861 x 14 $\rm arcsec^{2}$), allowing for better subtraction of night sky emission and scattered moonlight as compared to the B5-decker, was used for most of the subsequent observations.
We investigated the quality of the B5 RV set by fitting a 2-Keplerian model to subsets of the data, analyzing the best-fit parameters. The importance of the consistency of the extracted semi-amplitude in time for subsets of data is discussed in Hara et al. (2022a). For this analysis, we used PyORBIT (Malavolta 2016), fixing the periods and transit times of planets b and c and assuming circular orbits. The only free parameters were the semi-amplitudes of the two planets and the stellar RV jitter. Since planet b has a very short orbital period, we expect to get consistent semi-amplitudes for this planet across the data as long as we include a sufficient number of RVs. Fitting this model to the first 17 B5-decker observations in W16, we measured a remarkably large semi-amplitude of 4.8 m s-1for planet b followed by a drop to 1.6 m s-1for the closest 2-month chunk of C2-decker RVs. We also compared the B5-decker RVs with RV chunks of similar duration within the HARPS-N data with and without the YARARA correction. For this, we randomly selected 2-month windows of data with at least 15 observations, to match the characteristics of the B5 RV set, and fitted the 2-Keplerian model described above. The median number of included measurements was equal to 20, which is comparable to the number of B5-decker observations. The extracted semi-amplitude of planet b reached around 4 m s-1for only very few subsets of subsets of data, none reached 4.8 m s-1as extracted for the B5-decker RVs, and we found no semi-amplitude jumps as large as the one between the B5 and the C2-decker RVs. Therefore, the B5 data appear to be contaminated. Due to the B5 set consisting of few and noisy measurements, complicating a meaningful extraction of orbital information, and to keep the dataset uniform, we included only the C2 RVs.
As for the HARPS-N data, we searched for possible outliers in the HIRES C2-decker RVs with the Chauvenet’s criterion, and found four measurements at the epochs 2455314.006, 2455344.978, 2456908.977 and 2456909.885 $\rm BJD_{UTC}$, which were thus discarded.
@@ 3.1 Transit timing variations and parameters of Kepler-10 c
To compute the Kepler-10 c transit timing variations, we first fitted each transit independently by i) using the same differential evolution Markov chain Monte Carlo (DE-MCMC; Ter Braak 2006; Eastman et al. 2013) Bayesian framework as in D14; ii) considering intervals of the light curve centred at the predicted times of the Kepler-10 c transits from the linear ephemeris in D14, with a temporal window of twice the transit duration; iii) adopting a 2-planet transit model in case of superposition of the Kepler-10 b and Kepler-10 c transits, by fixing the parameters of the Kepler-10 b transits to the solution of D14; iv) imposing a Gaussian prior on the stellar density, as derived from previous asteroseismic analyses of the Kepler light curve; and v) fixing both the orbital period of planet c and the coefficients of the quadratic limb-darkening law to those found by D14 (no significant changes were noted when adopting instead Gaussian priors with the values and uncertainties given in D14). For the first two transits observed in long-cadence mode (29.4 min), we oversampled the model to 1-min samples (Kipping 2010). The Kepler-10 c transit times are given in Table 2; their variations with respect to a linear ephemeris are shown in Fig. 3, and are fully consistent with those reported by both Kipping et al. (2015) and W16.
We performed a subsequent, simultaneous DE-MCMC modeling of all the Kepler-10 c transits as in D14, after adjusting the offsets from the previously computed O-C. The determined transit parameters are in excellent agreement with those from previous literature works, and are reported in Table 5.
@@
Figure 3: Kepler-10 c Observed-Calculated (O-C) diagram showing the TTV pattern. The calculated times ($T_{c,\mathrm{lin}}$) are computed from a linear ephemeris: $T_{c,\mathrm{lin}}=T_{\mathrm{ref}}+N\times P=2454971.678363\pm 0.000659+N\times 45.294278\pm 0.000039$, where $N$ is an integer number identifying each transit time with respect to the reference time $T_{\mathrm{ref}}$.
3.1.1 Analysis of the TTV signal
@@
We analyzed the TTV signal presented in Fig. 3 to see what constraints we can obtain on the presence of non-transiting planets, and to make a comparison with the signal found in the RV analysis. A periodogram of the TTVs shown in Fig. 3 yields a peak at $P_{\rm TTV}=466.2_{-16.2}^{+17.2}$ days, which encompasses with $1\sigma$ the value given by W16. TTVs with timescales of tens to a few hundred times the orbital period of the planets are typically attributed to proximity to a mean motion resonance (MMR) with an additional planet in the system, defined by $P_{\rm out}/P_{\rm in}=(k+q)/k$, with $k$ and $q$ integers. Either the pair is inside the mean motion resonance, in which case the period of the TTVs scales as $P\leavevmode\nobreak\ (M_{\rm p}/M_{\star})^{-2/3}$ (e.g., Nesvorný & Vokrouhlický 2016), or the pair of planets is relatively close to the mean motion resonance, in which case the main TTVs period depends only on the orbital periods of the planets (e.g., Lithwick et al. 2012). Furthermore, short-term TTVs can be attributed to synodic ‘chopping’ effects (e.g., Deck & Agol 2015).
@@
First, we assumed that the perturbing planet is inside a mean motion resonance with Kepler-10 c. Given its orbital period of 45.2943 d, we estimated the mass that a perturbing planet should have to generate TTVs with a period of $\sim 466$ days using the formulas from Nesvorný & Vokrouhlický (2016). We found that the perturbing planet should have a mass ranging between about one Saturn mass (for a 7:6 resonance, either inside or outside Kepler-10 c) and about 10 Jupiter masses (for a 2:1 resonance either inside or outside Kepler-10 c). In addition to the fact that such a massive object should have easily been found in the radial velocity data, these configurations are also at the edge of instability (Deck et al. 2013).
@@
We then assumed that the perturbing planet is not inside, but relatively close to, a MMR of the form $P_{\rm out}/P_{\rm in}=(k+q)/k$, with $q=1$ or $2$. The main TTV frequency now depends only on the orbital period of Kepler-10 c, which we fixed to 45.2943 d, and the perturbing planet, with a period called the “super period” $1/(k/P_{\rm in}-(k+q)/P_{\rm out})$ (e.g., Lithwick et al. 2012). For each resonance considered, the perturbing planet can be either the inner or the outer planet. Then, for each of these configurations, the perturbing planet can be on either side of the MMR, since we only measure the absolute value of the super period. Supposing that $P_{\rm TTV}$ is such a super period, we can derive the orbital period of the companion, depending on the MMR to which it is close. These potential companion periods are summarized in Table 3.
@@
Interestingly, the signal at $151.06\pm 0.48$ d found in the HARPS-N RVs (see Sects. 3.2, 3.3, and Table LABEL:table_harpsn_param) is $1\sigma$ consistent with the orbital period of $150.50_{-0.57}^{+0.59}$ d required to generate the observed TTV period if the companion is relatively close to the 3:1 MMR with Kepler-10 c on an outer orbit. Given the relatively low uncertainty on both these periods, the probability that this is due to chance is quite low. The TTV signal therefore agrees with an orbital period of $\sim 151$ days for Kepler-10 d. In addition, the above-mentioned 3:1 resonance is of second order ($q=2$), and so the amplitude of the TTVs induced by the proximity to this resonance vanishes for zero eccentricities Hadden & Lithwick (2016). This explains the relatively low TTV amplitude observed, but also implies that other TTV harmonics than the proximity to the 3:1 MMR might have comparable amplitudes.
We therefore verified if the TTV signal contains additional information on the interactions between planets c and d. Generalizing the results of (Deck & Agol 2015), we know that the TTV signal expands as function of the harmonics of the mean longitude of the perturbing planet ($\lambda_{\rm d}$). We show the TTVs folded with respect to these harmonics in Fig. 4. The amplitude of each harmonic $j\lambda_{\rm d}$ is the sum of the contribution of the synodic (i.e. conjunction) harmonic $j(\lambda_{\rm c}-\lambda_{\rm d})$ and the effect of the mean motion resonances of the form $P_{\rm d}/P_{\rm c}=j/m$, with $m$ an integer.
@@
As can be seen in Fig. 4, only the 3rd harmonic shows a significant TTV amplitude. Since it is this harmonics that contains the effect of the 3:1 MMR, we therefore conclude that the proximity to this resonance is likely dominating the TTV signal.
From this analysis we can draw three conclusions: i) the main periodicity observed in the TTVs of planet c can be explained by the signal at 151 d also seen in the HARPS-N RVs; ii) since the relative proximity to this resonance generates TTVs only for eccentric planets, we expect Kepler-10 c and d to have non-zero eccentricities; and iii) even low S/N TTVs can help to confirm periodic signals found in RVs.
@@
3.2.1 Generalized Lomb-Scargle periodograms and iterative planet fitting.
The peaks at the periods of Kepler-10 b and Kepler-10 c are clearly visible in the GLS periodogram of the HARPS-N RVs (Figs. 13 and 2). To search for additional signals, we first modeled the RV signals of Kepler-10 b and Kepler-10 c as in Sect. 3.3.1, removed them from the original RV time series, and ran again the GLS periodogram on the residuals of the 2-planet model. This procedure leads to a better fit of the planet signals than the sinusoidal fit from the GLS parameters, allowing in particular for a slightly eccentric Keplerian fit for Kepler-10 c. The periodogram of the residuals showed a third signal with a period of $150.8\pm 0.7$ d and theoretical False Alarm Probability (FAP) of $3.4\cdot 10^{-7}$ (see Fig. 13). We then performed a 3-planet modeling, searched for other signals with the GLS periodogram in the best-fit residuals, and found a fourth signal with a period of $83.08\pm 0.27$ d and FAP of $3.6\cdot 10^{-4}$ (Fig. 13); the power at $\sim 26$ d seen in Fig. 2 instead disappears, being somehow related to both the signals of the 3 planets and the data sampling. However, the inclusion of this fourth signal is not favored by Bayesian model comparisons (Sect. 3.3), and thus it will not be included in the final modeling. None of these signals was found in the usable stellar magnetic activity indicators, such as the S-index and the bisector of the cross-correlation function.
@@ 3.2.2 False Inclusion Probability
We first applied the Bayesian methodology described in Hara et al. (2022b), whose end goal is to compute directly the posterior probability of the presence of a planet. We denote by $T_{\mathrm{obs}}$ the total observation timespan, and define $\Delta\omega=1/T_{\mathrm{obs}}$. We considered a grid of tightly spaced frequencies $\omega_{k}$ spanning from 0 to 1.5 cycles per day; for each $\omega_{k}$ in the grid, we computed the true inclusion probability (TIP), defined as the posterior probability that there is a planet with an orbital frequency in $[\omega_{k}-\Delta\omega,\omega_{k}+\Delta\omega]$. As shown in Hara et al. (2024), this detection criterion is optimal in the sense that, provided the likelihood and prior models are correct, it maximizes true detections for a given tolerance to false ones.
The computation of the TIP was done for a certain choice of priors and likelihood functions. We employed uniform priors on angles, a Beta prior on eccentricity as in Kipping (2014), and a log-uniform prior on the period. As shown in Hara et al. (2022b), the prior on the semi-amplitude and noise models can lead to drastic differences in statistical significance. We used either a log-uniform prior on semi-amplitude, or a mixture Gaussian model on the mass, and a white noise model. The results are shown in Fig. 5.
In all cases, we obtained very significant signals at 0.83 and 44.9 days with a false inclusion probability (FIP, equal to 1 - TIP), that is a probability of not being present, of less than $10^{-12}$. Depending on the priors, the FIP of the 150-152 day signal varied from $7\cdot 10^{-2}$ to $5\cdot 10^{-6}$. We took a further step and computed the TIP marginalised on the two noise models, white and red (see B.1 for details). To further investigate the origin of the 151 days signal and try to evaluate whether it is strictly periodic or not, we applied the methods of Hara et al. (2022a). The analysis, presented in detail in Appendix B, shows that the 151 day signal is compatible with a purely sinusoidal signal.
@@
Figure 5: FIP periodograms of the Kepler-10 HARPS-N data processed with YARARA for different priors on semi amplitude. (a) is obtained with a log-uniform prior on semi-amplitude, (b) with a Gaussian mixture model on mass. In yellow we represent the true inclusion probability (TIP), the probability of having a planet in the range $[\omega-\Delta\omega,\omega+\Delta\omega]$ as a function of $\omega$. In blue, we represent $-\log_{10}FIP$, where FIP = 1 - TIP is the false inclusion probability.
3.3 Radial-velocity modeling

### v2

@@ 2.2.2 HIRES radial velocities
The HIRES Kepler-10 spectra were recorded using different deckers that define the length of the effective slit. The first 17 HIRES observations that spanned approximately 38 days were made with the B5-decker (0.861 x 3.5 $\rm arcsec^{2}$) [^2] https://www2.keck.hawaii.edu/inst/hires/manual2.pdf. The longer C2-decker (0.861 x 14 $\rm arcsec^{2}$), which allows for better subtraction of night sky emission and scattered moonlight as compared to the B5-decker, was used for most of the subsequent observations.
We investigated the quality of the B5 RV set by fitting a two-Keplerian (i.e., two-planet) model to subsets of the data and analyzing the best-fit parameters. The importance of the consistency of the extracted semiamplitude in time for subsets of data is discussed in Hara et al. (2022a). For this analysis, we used PyORBIT (Malavolta 2016) by fixing the periods and transit times of planets b and c, and assuming circular orbits. The only free parameters were the semiamplitudes of the two planets and the stellar RV jitter. Since planet b has a very short orbital period, we expect to get consistent semiamplitudes for this planet across the data as long as we include a sufficient number of RVs. Fitting this model to the first 17 B5-decker observations in W16, we measured a remarkably large semiamplitude of 4.8 m s-1for planet b followed by a drop to 1.6 m s-1for the closest two-month chunk of C2-decker RVs. We also compared the B5-decker RVs with RV chunks of similar duration within the HARPS-N data with and without the YARARA-v2 correction. For this, we randomly selected two-month windows of data with at least 15 observations, to match the characteristics of the B5 RV set, and fitted the two-Keplerian model described above. The median number of included measurements was equal to 20, which is comparable to the number of B5-decker observations. The extracted semiamplitude of planet b reached around 4 m s-1for only a very few subsets of the data; none reached 4.8 m s-1as extracted for the B5-decker RVs, and we found no semiamplitude jumps as large as the one between the B5 and the C2-decker RVs. Therefore, the B5 data appear to be contaminated. To avoid the few noisy B5 measurements, which complicate a meaningful extraction of orbital information, and keep the dataset uniform, we considered only the C2 RVs.
As for the HARPS-N data, we searched for possible outliers in the HIRES C2-decker RVs with Chauvenet’s criterion, and found four measurements at the epochs 2455314.006, 2455344.978, 2456908.977, and 2456909.885 $\rm BJD_{UTC}$, which were thus discarded.
@@ 3.1 Transit timing variations and parameters of Kepler-10 c
To compute the Kepler-10 c transit timing variations, we first fitted each transit independently by (i) using the same differential evolution Markov chain Monte Carlo (DE-MCMC; Ter Braak 2006; Eastman et al. 2013) Bayesian framework as in D14; (ii) considering intervals of the light curve centered at the predicted times of the Kepler-10 c transits from the linear ephemeris in D14, with a temporal window of twice the transit duration; (iii) adopting a two-planet transit model in case of superposition of the Kepler-10 b and Kepler-10 c transits, by fixing the parameters of the Kepler-10 b transits to the solution of D14; (iv) imposing a Gaussian prior on the stellar density, as derived from previous asteroseismic analyses of the Kepler light curve; and (v) fixing both the orbital period of planet c and the coefficients of the quadratic limb-darkening law to those found by D14 (no significant changes were noted when adopting instead Gaussian priors with the values and uncertainties given in D14). For the first two transits observed in long-cadence mode (29.4 min), we oversampled the model to 1-min samples (Kipping 2010). The Kepler-10 c transit times are given in Table 2; their variations with respect to a linear ephemeris are shown in Fig. 3, and are fully consistent with those reported by both Kipping et al. (2015) and W16.
We performed a subsequent, simultaneous DE-MCMC modeling of all the Kepler-10 c transits as in D14, after adjusting the observed-calculated (O-C) offsets from the previously computed transit times. The determined transit parameters are in excellent agreement with those from previous works, and are reported in Table 5.
@@
Figure 3: Kepler-10 c observed-calculated (O-C) diagram showing the TTV pattern. The calculated times ($T_{c,\mathrm{lin}}$) are computed from a linear ephemeris: $T_{c,\mathrm{lin}}=T_{\mathrm{ref}}+N\times P=2454971.678363\pm 0.000659+N\times 45.294278\pm 0.000039$, where $N$ is an integer number that identifies each transit time with respect to the reference time $T_{\mathrm{ref}}$.
3.2 Analysis of the transit timing variation signal
@@
We analyzed the TTV signal presented in Fig. 3 to see what constraints we can obtain on the presence of non-transiting planets, and to make a comparison with the signal found in the RV analysis. A periodogram of the TTVs shown in Fig. 3 yields a peak at $P_{\rm TTV}=466.2_{-16.2}^{+17.2}$ days, which encompasses with $1\sigma$ the value given by W16. TTVs with timescales of tens to a few hundred times the orbital period of the planets are typically attributed to proximity to a mean motion resonance (MMR) with an additional planet in the system, defined by $P_{\rm out}/P_{\rm in}=(k+q)/k$, with $k$ and $q$ integers. Either the pair is inside the MMR, in which case the period of the TTVs scales as $P\leavevmode\nobreak\ (M_{\rm p}/M_{\star})^{-2/3}$ (e.g., Nesvorný & Vokrouhlický 2016), or the pair of planets is relatively close to the MMR, in which case the main TTV periodicity depends only on the orbital periods of the planets (e.g., Lithwick et al. 2012). Furthermore, short-term TTVs can be attributed to synodic ”chopping” effects (e.g., Deck & Agol 2015).
@@
First, we assumed that the perturbing planet is inside a MMR with Kepler-10 c. Given the Kepler-10 c orbital period of 45.2943 d, we estimated the mass that a perturbing planet would need to generate TTVs with a period of $\sim 466$ days using the formulas from Nesvorný & Vokrouhlický (2016). We found that the perturbing planet should have a mass ranging between about one Saturn mass (for a 7:6 resonance, either inside or outside Kepler-10 c) and about 10 Jupiter masses (for a 2:1 resonance either inside or outside Kepler-10 c). In addition to the fact that such a massive object should have easily been found in the RV data, these configurations are also at the edge of instability (Deck et al. 2013).
@@
We then assumed that the perturbing planet is not inside, but relatively close to, a MMR of the form $P_{\rm out}/P_{\rm in}=(k+q)/k$, with $q=1$ or $2$. The main TTV frequency now depends only on the orbital period of Kepler-10 c, which we fixed to 45.2943 d, and the perturbing planet, with a period called the “super period” $1/(k/P_{\rm in}-(k+q)/P_{\rm out})$ (e.g., Lithwick et al. 2012). For each resonance considered, the perturbing planet can be either the inner or the outer planet. Then, for each of these configurations, the perturbing planet can be on either side of the MMR, since we only measure the absolute value of the super period. Supposing that $P_{\rm TTV}$ is such a super period, we can derive the orbital period of the companion, depending on the MMR it is close to. These potential companion periods are summarized in Table 3.
@@
Interestingly, the signal at $151.06\pm 0.48$ d found in the HARPS-N RVs (see Sects. 3.3, 3.4, and Table LABEL:table_harpsn_param) is $1\sigma$ consistent with the orbital period of $150.50_{-0.57}^{+0.59}$ d required to generate the observed TTV period if the companion is relatively close to the 3:1 MMR with Kepler-10 c on an outer orbit. Given the relatively low uncertainty on both these periods, the probability that this is due to chance is quite low. The TTV signal therefore agrees with an orbital period of $\sim 151$ d for Kepler-10 d. In addition, the abovementioned 3:1 resonance is of second order ($q=2$), and so the amplitude of the TTVs induced by the proximity to this resonance vanishes for zero eccentricities (Hadden & Lithwick 2016). This explains the relatively low TTV amplitude observed, but also implies that other TTV harmonics than the proximity to the 3:1 MMR might have comparable amplitudes.
We therefore verified if the TTV signal contains additional information on the interactions between planets c and d. Generalizing the results of Deck & Agol (2015), we know that the TTV signal expands as a function of the harmonics of the mean longitude of the perturbing planet ($\lambda_{\rm d}$). We show the TTVs folded with respect to these harmonics in Fig. 4. The amplitude of each harmonic $j\lambda_{\rm d}$ is the sum of the contribution of the synodic (i.e., conjunction) harmonic $j(\lambda_{\rm c}-\lambda_{\rm d})$ and the effect of the MMRs of the form $P_{\rm d}/P_{\rm c}=j/m$, with $m$ an integer. As can be seen in Fig. 4, only the third harmonic shows a significant TTV amplitude. Since it is this harmonic that contains the effect of the 3:1 MMR, we therefore conclude that the proximity to this resonance is likely dominating the TTV signal.
@@
From this analysis, we can draw three conclusions: (i) the main periodicity observed in the TTVs of planet c can be explained by the signal at 151 d also seen in the HARPS-N RVs; (ii) since the relative proximity to this resonance generates TTVs for eccentric planets only, we expect Kepler-10 c and d to have non-zero eccentricities; and (iii) even low S/N TTVs can help confirm periodic signals found in RVs.
@@
3.3.1 Generalized Lomb-Scargle periodograms and iterative planet fitting
The peaks at the periods of Kepler-10 b and Kepler-10 c are clearly visible in the GLS periodogram of the HARPS-N RVs (Figs. 12 and 2). To search for additional signals, we first modeled the RV signals of Kepler-10 b and Kepler-10 c as in Sect. 3.4.1, removed them from the original RV time series, and ran again the GLS periodogram on the residuals of the two-planet model. This procedure leads to a better fit of the planet signals than the sinusoidal fit from the GLS parameters, allowing in particular a slightly eccentric Keplerian fit for Kepler-10 c. The periodogram of the residuals showed a third signal with a period of $150.8\pm 0.7$ d and a theoretical false alarm probability (FAP) of $3.4\cdot 10^{-7}$ (see Fig. 12). We then performed a three-planet modeling, searched for other signals with the GLS periodogram in the best-fit residuals, and found a fourth signal with a period of $83.08\pm 0.27$ d and a FAP of $3.6\cdot 10^{-4}$ (Fig. 12); the power at $\sim 26$ d seen in Fig. 2 instead disappeared, being somehow related to both the signals of the three planets and the data sampling. However, the inclusion of this fourth signal was not favored by Bayesian model comparisons (Sect. 3.4), and thus it was not considered in the final modeling. None of these signals was found in the usable stellar magnetic activity indicators, such as the S-index and the bisector of the CCF.
@@ 3.3.2 False inclusion probability
Figure 5: False inclusion probability periodograms of the Kepler-10 HARPS-N data processed with YARARA-v2 for different priors on semiamplitude. (a) was obtained with a log-uniform prior on semiamplitude, (b) with a Gaussian mixture model on mass. In yellow we represent the true inclusion probability (TIP), the probability of having a planet in the range $[\omega-\Delta\omega,\omega+\Delta\omega]$ as a function of $\omega$. In blue, we represent $-\log_{10}FIP$, where FIP = 1 - TIP is the false inclusion probability.
We first applied the Bayesian methodology described in Hara et al. (2022b), whose end goal is to compute directly the posterior probability of the presence of a planet. By $T_{\mathrm{obs}}$ we denote the total observation timespan, and define $\Delta\omega=1/T_{\mathrm{obs}}$. We considered a grid of tightly spaced frequencies $\omega_{k}$ spanning from 0 to 1.5 cycles per day; for each $\omega_{k}$ in the grid, we computed the true inclusion probability (TIP), defined as the posterior probability that there is a planet with an orbital frequency in $[\omega_{k}-\Delta\omega,\omega_{k}+\Delta\omega]$. As shown in Hara et al. (2024), this detection criterion is optimal in the sense that, provided the likelihood and prior models are correct, it maximizes true detections for a given tolerance to false ones.
The computation of the TIP was done for a certain choice of priors and likelihood functions. We employed uniform priors on angles, a Beta prior on eccentricity as in Kipping (2014), and a log-uniform prior on the period. As shown in Hara et al. (2022b), the prior on the semiamplitude and noise models can lead to drastic differences in statistical significance. We used either a log-uniform prior on semiamplitude, or a mixture Gaussian model on the mass, and a white noise model. The results are shown in Fig. 5.
In all cases, we obtained very significant signals at 0.83 and 44.9 d with a false inclusion probability (FIP, equal to 1 - TIP), that is, a probability of not being present, of less than $10^{-12}$. Depending on the priors, the FIP of the 150-152 d signal varied from $7\cdot 10^{-2}$ to $5\cdot 10^{-6}$. We took a further step and computed the TIP marginalized on the two noise models, white and red (see B.1 for details). To further investigate the origin of the 151 d signal and try to evaluate whether it was strictly periodic or not, we applied the methods of Hara et al. (2022a). The analysis presented in detail in Appendix B shows that the 151 d signal is compatible with a purely sinusoidal signal.
3.4 Modeling of radial velocities